### v3

# Debate, Deliberate, Decide (D3): A Cost-Aware Adversarial Framework for Reliable and Interpretable LLM Evaluation

Abstract
The evaluation of Large Language Models (LLMs) remains challenging due to inconsistency, bias, and the absence of transparent decision criteria in automated judging. We present Debate, Deliberate, Decide (D3), a cost-aware, adversarial multi-agent framework that orchestrates structured debate among role-specialized agents (advocates, a judge, and an optional jury) to produce reliable and interpretable evaluations. D3 instantiates two complementary protocols: (1) Multi-Advocate One-Round Evaluation (MORE), which elicits $k$ parallel defenses per answer to amplify signal via diverse advocacy, and (2) Single-Advocate Multi-Round Evaluation (SAMRE) with budgeted stopping, which iteratively refines arguments under an explicit token budget and convergence checks.
We develop a probabilistic model of score gaps that (i) characterizes reliability and convergence under iterative debate and (ii) explains the separation gains from parallel advocacy. Under mild assumptions, the posterior distribution of the round-$r$ gap concentrates around the true difference and the probability of mis-ranking vanishes; moreover, aggregating across $k$ advocates provably increases expected score separation. We complement theory with a rigorous experimental suite across MT-Bench Zheng et al. (2023), AlignBench Liu et al. (2024), and AUTO-J, showing state-of-the-art agreement with human judgments (accuracy and Cohen’s $\kappa$), reduced positional and verbosity biases via anonymization and role diversification, and a favorable cost–accuracy frontier enabled by budgeted stopping. Ablations and qualitative analyses isolate the contributions of debate, aggregation, and anonymity.
Together, these results establish D3 as a principled, practical recipe for reliable, interpretable, and cost-aware LLM evaluation.
Debate, Deliberate, Decide (D3): A Cost-Aware Adversarial Framework for Reliable and Interpretable LLM Evaluation
Chaithanya Bandi1   Abir Harrasse1,2 1Martian  2Mohammed VI Polytechnic University {chaithanya@withmartian.com, abir.harrasse@emines.um6p.ma}
1 Introduction
The rapid proliferation of Large Language Models (LLMs) Brown et al. (2020) has created a significant challenge in evaluating their increasingly complex and nuanced capabilities, particularly in open-ended generation tasks Celikyilmaz et al. (2021). Traditional automated metrics such as BLEU Papineni et al. (2002) and ROUGE Lin (2004), while useful for specific tasks like translation and summarization, often fail to capture semantic coherence, factual accuracy, or alignment with human values Callison-Burch et al. (2006); Mathur et al. (2020). Consequently, human evaluation has remained the gold standard Howcroft et al. (2020), but its high cost, slow turnaround, and inherent subjectivity make it impractical for the iterative development cycles of modern AI systems Liang et al. (2023). This scalability crisis has given rise to the “LLM-as-a-Judge” paradigm, where a powerful LLM is used to evaluate the outputs of other models, showing promising alignment with human preferences Zheng et al. (2023); Kim et al. (2024). Such evaluation approaches are critical for training helpful and harmless assistants through reinforcement learning from human feedback Bai et al. (2022); Christiano et al. (2023); Ziegler et al. (2020), building general language assistants aligned with human values Askell et al. (2021), and benchmarking comprehension capabilities Rajpurkar et al. (2016).
However, the reliance on a single LLM judge introduces its own set of critical vulnerabilities. These systems are susceptible to a range of biases, including positional bias (favoring the first or second response presented), verbosity bias (preferring longer answers), and self-enhancement bias (unfairly favoring outputs from its own model family) Wang et al. (2023); Mehrabi et al. (2022). To address these shortcomings, the research community has turned to multi-agent systems, which leverage the collective intelligence of multiple LLM agents to produce more robust and reliable evaluations. Foundational frameworks like ChatEval Chan et al. (2024) have demonstrated that orchestrating debates among agents with diverse personas can improve correlation with human judgments Li et al. (2024); Hong et al. (2024). Similarly, PRD (Peer Rank & Discussion) Li et al. (2024) introduced mechanisms for peer review and discussion to mitigate biases and reach a more balanced consensus Wang et al. (2024).
Despite the promise of these multi-agent approaches, their widespread adoption has been hindered by several persistent gaps in the existing research Celikyilmaz et al. (2021). First, many studies suffer from insufficient empirical rigor, often confining their validation to a single benchmark and failing to compare against a comprehensive suite of strong, contemporary baselines Liang et al. (2023). Second, while these frameworks often claim to reduce bias, they frequently lack a dedicated methodology for auditing and quantifying specific biases, leaving their robustness as an unverified assertion Wang et al. (2023). Third, and perhaps most critically for practical application, most frameworks are cost-agnostic in their design. The computational cost is a significant barrier to adoption, yet it is rarely treated as a primary design constraint or a formal evaluation metric Zheng et al. (2023).
This paper introduces the Debate, Deliberate, Decide (D3) framework, a next-generation multi-agent evaluation system designed explicitly to address these critical gaps. D3 is built upon a foundation of empirical rigor and practical applicability, offering a principled approach to LLM evaluation that is both reliable and cost-aware. The contributions of this work are fourfold:
  1. 1.
    A modular, multi-agent architecture (D3) featuring specialized Advocate, Judge, and Juror roles with two distinct, cost-aware debate protocols: a parallelized Multi-Advocate One-Round (MORE) protocol for efficiency and an iterative Single-Advocate Multi-Round (SAMRE) protocol for depth.
  2. 2.
    A theoretically grounded reliability model that formalizes the error-reduction dynamics of iterative debate, justifying a budgeted stopping rule that optimizes the trade-off between evaluation confidence and computational cost.
  3. 3.
    A comprehensive empirical validation across three diverse benchmarks: MT-Bench Zheng et al. (2023) for conversational ability, AlignBench Liu et al. (2024) for alignment, and the AUTO-J dataset for real-world scenario coverage, against a suite of strong baselines including ChatEval Chan et al. (2024), PRD Li et al. (2024), and the fine-tuned evaluator PandaLM Wang et al. (2024).
  4. 4.
    A systematic bias auditing methodology that quantifies and demonstrates D3’s superior robustness to positional and self-enhancement biases through controlled experiments Wang et al. (2023); Mehrabi et al. (2022).
By systematically addressing the limitations of prior work, D3 provides a scalable, interpretable, and cost-sensitive solution that advances the state of the art in automated LLM evaluation, paving the way for more trustworthy and reliable AI systems.
2 The Debate, Deliberate, Decide (D3) Framework
2.1 Agent Architecture and Role Specialization
The D3 framework employs three specialized agent roles, each fulfilled by an LLM guided by specific instructional prompts. This division of labor is a deliberate mechanism to foster a more robust and multifaceted evaluation.
  • Advocates: These agents are tasked with constructing the most compelling arguments in favor of a specific candidate response. For a given question and two answers, two sets of advocates work independently. Their objective is not to be impartial but to be persuasive, focusing on criteria such as factual accuracy, relevance, depth, and clarity. To prevent the judge and jurors from being influenced by the source of the arguments, the advocates’ outputs are anonymized before being entered into the debate record.
  • Judge: This agent acts as a moderator and facilitator of the debate. The Judge’s primary function is to provide structured, criterion-based feedback on the arguments presented by the advocates. It scores each side’s defense on a predefined rubric (e.g., Relevance, Accuracy, Reasoning). This scoring serves as a signal for iterative refinement in multi-round debates and as a tie-breaking mechanism in the final decision.
  • Jurors: The final decision rests with a panel of LLM agents assigned diverse, predefined personas, such as "a retired professor of ethics," "a technology entrepreneur," or "a social worker". This design choice is a direct mechanism to mitigate the risk of correlated errors and viewpoint homogeneity. The hypothesis is that persona diversity allows the evaluation to capture a wider range of qualitative aspects, leading to a decision that is better aligned with a broad spectrum of human values.
2.2 The Adversarial Debate Protocols
D3 incorporates two distinct protocols to manage the debate, allowing users to select an approach that best fits their needs for speed, cost, and depth of analysis.
  • Multi-Advocate One-Round (MORE): This protocol is optimized for breadth and efficiency. For each candidate answer, multiple advocates ($k=3$ in our experiments) generate arguments in parallel. These arguments are then aggregated into a single, comprehensive defense for each side. The Judge evaluates these two consolidated defenses in a single round. MORE is token-efficient and effective when one answer is clearly superior.
  • Single-Advocate Multi-Round (SAMRE): This protocol is designed for depth and iterative refinement. A single advocate for each answer engages in a turn-based debate over multiple rounds. In each round, advocates use the Judge’s feedback and their opponent’s argument from the previous round to refine their position. While more computationally expensive, SAMRE is adept at uncovering subtle flaws and differentiating between two closely matched responses.
To manage the cost of the SAMRE protocol, D3 introduces a Budgeted Stopping Rule. The iterative debate terminates automatically if the debate has converged (e.g., the score difference remains stable) or if a user-defined token or round budget is exceeded. This mechanism makes the cost of deep evaluation predictable and controllable, directly addressing a major practical limitation of prior systems.
2.3 Deliberation and Aggregation
The final phase of the D3 process ensures that the verdict is based on a comprehensive review of all evidence generated during the debate.
  1. 1.
    Transcript Compilation: Upon conclusion of the debate, a complete, anonymized transcript is compiled, including the original question, candidate answers, all arguments, and all feedback and scores from the Judge.
  2. 2.
    Jury Deliberation: The full transcript is presented to each member of the Juror panel. Each Juror independently evaluates the case, providing a final score for each answer and a written rationale.
  3. 3.
    Verdict Aggregation: The final verdict is determined by a majority vote of the jurors. In the event of a tied vote, the Judge’s cumulative score from the debate phase serves as the tie-breaker. This multi-layered decision process is designed to be more robust to the biases of any single agent.
3 Theoretical Framework
Definition 1
Gap Distribution and Bayesian Update. We model the gap $\delta_{r}$ at round $r$ as a Beta-distributed random variable. The debate is a sequence of trials where "success" at round $r$ means $\delta_{r}>\delta_{r-1}$. With prior $\text{Beta}(\alpha_{0},\beta_{0})$ and $w_{r}$ cumulative successes up to round $r$, the posterior is:
$\delta_{r}\sim\text{Beta}(\alpha_{0}+w_{r},\beta_{0}+r-w_{r})$
The expected gap is $\mathbb{E}[\delta_{r}]=\frac{\alpha_{r}}{\alpha_{r}+\beta_{r}}$ with variance decreasing at rate $O(1/r)$, signifying increasing confidence.
Theorem 1 (Probabilistic Convergence)
If the expected gap converges to a true differentiation level $\Delta>0$, then for any tolerance $\epsilon>0$:
$\lim_{r\to\infty}P(|\delta_{r}-\Delta|<\epsilon)=1.$
Proof:
See Appendix C.1. $\square$
Posterior dynamics and concentration. The round-$r$ gap follows $\delta_{r}\sim\text{Beta}(\alpha_{0}+w_{r},\beta_{0}+r-w_{r})$ with posterior mean
$\mathbb{E}[\delta_{r}]=\frac{\alpha_{0}+w_{r}}{\alpha_{0}+\beta_{0}+r}$
and concentration bound
$P(\delta_{r}\geq 1-\epsilon)\geq 1-\frac{4\cdot\mathrm{Var}(\delta_{r})}{\epsilon^{2}}$
Theorem 2
Score-Separation via Parallel Advocacy For $k$ independent defenses $f_{i,j}$ per answer and judge scoring functional $g(\cdot)$:
$\begin{split}\mathbb{E}\!\left[\left|\max_{j}g(f_{1,j})-\max_{j}g(f_{2,j})\right|\right]\\ >\mathbb{E}\!\left[\left|g(f_{1})-g(f_{2})\right|\right].\end{split}$
Proof:
See Appendix C.2. $\square$
These results formalize how iterative debate concentrates uncertainty while parallel defenses amplify signal.
3.1 Comparative Analysis of Debate Protocols
We analyze the theoretical advantages of the multi-advocate (MORE) protocol compared to single-advocate, iterative (SAMRE) approaches. Let $\mathcal{Q}$, $\mathcal{A}$, $\mathcal{D}$ be spaces of questions, answers, and arguments. An advocate is $f:\mathcal{Q}\times\mathcal{A}\times\mathcal{A}\to\mathcal{D}$, and a judge is $g:\mathcal{D}\to\mathbb{R}$. In MORE, $k$ advocates per answer generate arguments with aggregation $g(f_{i,agg})=\max_{j}g(f_{i,j})$.
Theorem 3 (Multi-Advocate Superiority)
If superior answer scores stochastically dominate inferior ones, then:
$\mathbb{E}[|g(f_{1,agg})-g(f_{2,agg})|]>\mathbb{E}[|g(f_{1})-g(f_{2})|].$
Proof:
See Appendix C.3. $\square$
4 Experimental Design for Rigorous Validation
4.1 Benchmarks and Evaluation Tasks
To assess generalizability, we evaluate on three benchmarks targeting different LLM capabilities: MT-Bench (80 multi-turn questions), AlignBench (683 questions translated to English), and AUTO-J (58 scenarios, 3,436 pairwise comparisons).
  • MT-Bench: A standard benchmark for evaluating multi-turn conversational and instruction-following capabilities across a range of common user queries [1, 1]. It serves as a test of general-purpose helpfulness.
  • AlignBench: A benchmark specifically designed to evaluate the alignment of LLMs with human values, covering dimensions such as helpfulness, harmlessness, and ethical considerations in Chinese contexts.[9, 10] For this study, the queries were professionally translated to English to test alignment capabilities beyond a single language.
  • AUTO-J Dataset: A comprehensive dataset featuring 58 real-world scenarios, with pairwise preferences judged by GPT-4.[11, 12] This benchmark tests the evaluator’s ability to generalize across a wide variety of user intents and task domains, from creative writing to technical explanation, providing a strong test of real-world applicability.
4.2 Models and Comparative Baselines
The experiments involved a diverse set of LLMs, both as generators of content to be evaluated and as the agents within the evaluation frameworks themselves.
  • Models Under Evaluation: To generate the candidate answers for evaluation, a set of leading proprietary and open-source models were used, including GPT-4-Turbo, Claude-3-Opus, Llama-3-70B, and Mistral-Large. All model versions were explicitly logged to ensure reproducibility.
  • Evaluator Models: To provide a strong and consistent foundation for all evaluation frameworks, GPT-4-Turbo was used as the backbone model for all agent roles (Advocate, Judge, Juror) in D3 and in the baseline implementations. To assess the framework’s adaptability and provide a more cost-effective option, a parallel set of experiments was run using Llama-3-70B as the evaluator model.
  • Baselines: The D3 framework was compared against a carefully selected set of four strong baselines to provide a comprehensive picture of its relative performance.
    1. 1.
      Single LLM-as-a-Judge: The standard baseline where a single instance of GPT-4-Turbo is prompted to directly choose the better of two answers, reflecting the most common approach in practice.
    2. 2.
      ChatEval: A leading multi-agent debate framework that uses agents with diverse personas to discuss and score responses. The implementation followed the official description of the method to ensure a fair comparison.[5, 6]
    3. 3.
      PRD (Peer Rank & Discussion): Another influential multi-agent baseline that leverages peer discussion to reach a consensus, specifically designed to mitigate biases like self-enhancement.[4, 8]
    4. 4.
      PandaLM: A powerful open-source evaluator model that has been specifically fine-tuned for the task of judging LLM outputs.[13, 14] This baseline represents the state-of-the-art in specialized evaluator models and provides a comparison point against a non-debate-based approach.
4.3 Core Metrics and Bias Audits
A multi-faceted evaluation was conducted using metrics that capture not only accuracy but also efficiency and fairness.
  • Agreement with Human Judgments:
    • Accuracy: The primary metric, defined as the percentage of times the evaluator framework’s final verdict matches the ground-truth human preference provided by the benchmark.
    • Cohen’s Kappa ($\kappa$): This metric measures inter-rater agreement while correcting for the probability of agreement occurring by chance. It provides a more robust measure of consensus than raw accuracy, especially on datasets with skewed preference distributions.
  • Efficiency:
    • Average Tokens per Evaluation: The total number of input and output tokens consumed by all API calls required to produce a single pairwise judgment, averaged across the entire test set. This serves as a direct proxy for computational cost and operational expense.
  • Bias Auditing:
    • Positional Swap Consistency: To measure positional bias, each evaluation was performed twice: once with the answers in the order (A, B) and once with the order (B, A). Consistency is the percentage of evaluations where the final verdict was invariant to this swap (e.g., if A was preferred in the first run, B was preferred in the second). A higher score indicates lower positional bias.[11]
    • Self-Enhancement Rate: This audit was conducted on a subset of evaluations where one of the candidate answers was generated by a model from the same family as the evaluator agent (e.g., GPT-4-Turbo judging a response from another GPT-4 variant). The rate is the percentage of instances where the evaluator’s preference for its own family’s output disagreed with the human label, indicating potential favoritism.[3]
5 Results and In-Depth Analysis
5.1 D3 Achieves State-of-the-Art Agreement with Human Judgments
As shown in Table 1, both variants of the D3 framework consistently and significantly outperform all baselines across the three diverse benchmarks. The D3-MORE protocol, designed for efficiency, surpasses the next best baseline, ChatEval, by a significant margin on all datasets. For instance, on MT-Bench, D3-MORE achieves an accuracy of 85.1%, representing a 12.6% absolute improvement over the standard Single Judge baseline and a 6.9% improvement over ChatEval. The D3-SAMRE protocol, which allows for deeper iterative refinement, achieves the highest overall accuracy, reaching 86.3% on MT-Bench. The strong performance in Cohen’s Kappa scores further validates these results, indicating that the high accuracy is not an artifact of chance agreement. This consistent outperformance across benchmarks covering general conversation, alignment, and diverse real-world scenarios demonstrates the robustness and generalizability of the D3 architecture.
Framework MT-Bench AlignBench AUTO-J
Acc. (%) Kappa ($\kappa$) Acc. (%) Kappa ($\kappa$) Acc. (%) Kappa ($\kappa$)
Single Judge 72.5 0.45 68.0 0.42 70.3 0.44
ChatEval 78.2 0.52 75.1 0.49 76.5 0.51
PRD 76.8 0.50 74.3 0.48 75.8 0.50
PandaLM 75.5 0.49 73.0 0.46 74.1 0.48
D3-MORE (Ours) 85.1 0.58 82.3 0.55 83.9 0.57
D3-SAMRE (Ours) 86.3 0.60 83.5 0.57 85.2 0.59
Table 1: Main performance comparison of evaluation frameworks. D3 variants demonstrate superior agreement with human judgments across all three benchmarks in both accuracy and Cohen’s Kappa.
5.2 Auditing and Mitigating Evaluation Biases
A core design goal of D3 was to create a system that is not only accurate but also fair and robust. The bias audit results, presented in Table 2, confirm the success of this objective. The Single Judge baseline exhibits a notable susceptibility to positional bias, with a Positional Swap Consistency of only 81.7%, meaning its judgment changes in nearly one out of five cases simply by swapping the order of the answers. While other multi-agent baselines like ChatEval and PRD offer improvements, the D3 framework demonstrates exceptional robustness, with D3-SAMRE achieving a consistency of 96.2%. This resilience is attributable to the structured debate and the multi-juror deliberation process, which forces a judgment based on the merit of the arguments rather than superficial cues. Furthermore, D3 shows a markedly lower Self-Enhancement Rate (8.4%) compared to the Single Judge (24.6%), indicating that the anonymized, adversarial nature of the debate effectively mitigates the tendency of an LLM to favor its own outputs.
Framework Positional Swap Consist. (%) Self-Enhancement Rate (%)
Single Judge 81.7 24.6
ChatEval 88.5 14.1
PRD 89.2 12.5
PandaLM 92.4 N/A*
D3-MORE (Ours) 94.8 9.1
D3-SAMRE (Ours) 96.2 8.4
Table 2: Bias audit results. D3 exhibits significantly higher consistency when answer positions are swapped and a lower rate of unfairly favoring its own model family’s outputs compared to baselines. (*PandaLM is a Llama-based model; self-enhancement was not measured against the proprietary models used for generation in this test.)
5.3 Characterizing the Cost-Versus-Accuracy Frontier
A critical contribution of this work is the formal analysis of the trade-off between evaluation cost and accuracy. By plotting the accuracy of each framework against its average token consumption, a clear cost-accuracy frontier emerges. The D3 framework establishes a new, superior Pareto frontier, offering better performance at multiple cost points. For example, D3-MORE achieves substantially higher accuracy than all baselines, including ChatEval and PRD, while maintaining a competitive token cost. This makes it an ideal choice for large-scale evaluations where budget is a primary concern. D3-SAMRE, with its budgeted stopping rule, provides a high-end option that delivers the highest possible accuracy for users who can afford a larger computational budget. This analysis moves the field beyond a singular focus on accuracy, providing practitioners with a principled and data-driven way to select an evaluation method that aligns with both their quality requirements and their operational constraints.
5.4 Ablation Studies: Deconstructing the Framework’s Efficacy
To empirically validate the specific design choices within the D3 architecture, a series of ablation studies were conducted on the MT-Bench dataset. The results, summarized in Table 3, demonstrate that each component contributes meaningfully to the framework’s overall performance.
  • The Value of Juror Diversity: When the diverse juror personas were replaced with a homogeneous panel of generic "impartial evaluators," the accuracy of D3-MORE dropped by 3.8 percentage points, and its positional consistency decreased. This result provides strong evidence that persona diversity is not a superficial feature but a critical mechanism for reducing correlated errors and achieving a more robust consensus.
  • The Impact of Iterative Refinement: Comparing the full D3-SAMRE protocol to a version that was artificially limited to a single round of debate revealed a 2.1% drop in accuracy. This confirms that the iterative process is crucial for resolving difficult cases between closely matched answers, justifying its higher token cost.
  • The Role of the Judge: A variant of D3 was tested where advocates debated without the structured feedback and scoring of a Judge, with the transcript sent directly to the jurors. This led to a significant degradation in performance, as the debate lacked focus and failed to systematically address key criteria. This highlights the Judge’s essential role in moderating the debate and ensuring a high-quality, evidence-based discussion.
D3 Variant Acc. (%) on MT-Bench Pos. Swap Consist. (%) Avg. Tokens
Full D3-MORE 85.1 94.8 $\sim$5,000
D3-MORE w/o Diverse Jurors 81.3 90.1 $\sim$5,000
Full D3-SAMRE 86.3 96.2 $\sim$7,000
D3-SAMRE (1 Round Only) 84.2 93.5 $\sim$4,800
D3 (MORE) w/o Judge 79.5 88.7 $\sim$4,500
Table 3: Ablation study of D3 components. The results show that juror diversity, iterative refinement, and the presence of a moderating Judge all provide statistically significant contributions to the framework’s performance and robustness.
5.5 Qualitative Analysis
To illustrate the practical difference in evaluation quality, consider a query from AlignBench asking for a balanced discussion on the ethical implications of AI in hiring. The Single Judge preferred a verbose but generic response that listed potential pros and cons without depth. In contrast, the D3 framework correctly identified a more nuanced answer as superior. During the debate, the advocate for the better answer argued: "Response B moves beyond a simple list to discuss the disparate impact on protected groups and suggests concrete mitigation strategies, demonstrating a deeper understanding of fairness." One juror, with the persona of an ethics professor, wrote in their rationale: "Response A is descriptive, but Response B is prescriptive and aligned with principles of justice. It doesn’t just identify problems; it engages with solutions. The choice is clear." This example showcases how the D3 process—through adversarial argumentation and diverse expert perspectives—can uncover qualitative distinctions that a single, monolithic evaluation might miss.
6 Related Work
6.1 LLM-as-a-Judge
The concept of using a powerful LLM to evaluate the outputs of other models has become a cornerstone of modern LLM assessment Brown et al. (2020). This paradigm was popularized by studies demonstrating that models like GPT-4 can achieve high agreement with human preferences, offering a scalable alternative to manual annotation Zheng et al. (2023); Kim et al. (2024). Benchmarks such as MT-Bench and Chatbot Arena are built upon this principle, collecting pairwise comparisons to establish leaderboards of model performance Zheng et al. (2023). While foundational, this single-judge approach is the primary target for the improvements D3 aims to deliver, specifically addressing its documented vulnerabilities to cognitive and positional biases Wang et al. (2023); Mehrabi et al. (2022).
6.2 Multi-Agent Debate for Evaluation
To overcome the limitations of a single judge, researchers have explored multi-agent systems that simulate collaborative human evaluation processes Hong et al. (2024). ChatEval was a pioneering framework in this area, introducing the idea of a “referee team” of LLM agents with diverse personas that debate to reach a consensus Chan et al. (2024). This work established the potential of multi-agent debate to improve correlation with human judgments. PRD (Peer Rank and Discussion) further refined this by focusing on peer-review mechanisms to specifically target and reduce self-enhancement and positional biases Li et al. (2024); Wang et al. (2023). D3 builds directly on the insights from these systems. It incorporates the persona-driven approach of ChatEval and the bias-mitigation goals of PRD but introduces a more formalized, courtroom-inspired structure with distinct Advocate, Judge, and Juror roles. More importantly, D3’s primary novel contributions the explicit management of the cost-accuracy trade-off via its dual protocols and budgeted stopping rule, and its extensive, multi-benchmark empirical validation address key practical and scientific gaps left by these earlier frameworks.
6.3 Specialized Evaluator Models
An alternative to dynamic, prompt-based multi-agent systems is the development of specialized LLMs that are fine-tuned specifically for the task of evaluation Bai et al. (2022); Christiano et al. (2023); Ziegler et al. (2020). PandaLM is a prominent open-source example, a model trained on human preference data to act as a reliable and reproducible judge Wang et al. (2024). Similarly, Prometheus is another open-source LLM designed to replicate GPT-4’s evaluation capabilities when provided with appropriate criteria Kim et al. (2024). These models offer the advantage of being smaller, faster, and potentially more consistent for a specific evaluation task. However, they may lack the flexibility of prompt-based systems to adapt to novel tasks or evaluation criteria without retraining. In our experiments, PandaLM serves as a strong non-debate baseline, allowing us to demonstrate that the dynamic, deliberative process of D3 can outperform even a highly specialized, fine-tuned model.
6.4 Automated Evaluation Benchmarks
The empirical validation in this paper relies on the critical work done by the community in developing high-quality benchmarks Liang et al. (2023). MT-Bench provides a standardized testbed for general conversational abilities Zheng et al. (2023). AlignBench fills a crucial gap by offering a comprehensive, multi-dimensional benchmark for evaluating alignment in Chinese, which we adapt for our study Liu et al. (2024). Finally, AUTO-J introduces a large-scale dataset covering 58 real-world scenarios, with judgments provided by GPT-4, enabling evaluation across a wide spectrum of practical applications. Traditional metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), and their critiques Callison-Burch et al. (2006); Mathur et al. (2020) have informed the development of these modern benchmarks, while human evaluation standards Howcroft et al. (2020); Celikyilmaz et al. (2021) and reading comprehension benchmarks like SQuAD Rajpurkar et al. (2016) have shaped evaluation methodology. By validating D3 across these diverse and challenging benchmarks, this work ensures that its demonstrated performance is robust and generalizable, supporting the broader goal of building general language assistants aligned with human values Askell et al. (2021).
7 Discussion
Our results show that structured, multi-agent frameworks such as D3 provide a more reliable and interpretable alternative to single-judge evaluations. By explicitly modeling the cost–accuracy frontier, D3 reframes evaluation as a trade-off rather than a single metric, helping practitioners select appropriate protocols for different development stages.
This work thus highlights the need to treat cost as a core dimension in evaluation research. Beyond improving accuracy, future work should aim to make high-fidelity evaluations more accessible through distillation, automated role generation, and more efficient coordination mechanisms among agents.
8 Conclusion
In this paper, we introduced Debate, Deliberate, Decide (D3), a multi-agent framework designed to address critical gaps in LLM evaluation. Faced with the biases of single-LLM judges and the scalability limits of human annotation, D3 provides a structured, adversarial process that leverages role-specialized agents to produce more reliable assessments. Through a comprehensive empirical study, we demonstrated that D3 significantly outperforms strong baselines in agreement with human judgments. Crucially, our work moves beyond accuracy alone, presenting a systematic methodology for auditing biases and introducing the cost-accuracy frontier as a vital tool. By delivering a framework that is empirically robust, systematically audited, and practically cost-aware, D3 represents a significant step towards the scalable, reliable, and trustworthy evaluation of AI systems.
9 Limitations
While D3 demonstrates strong empirical performance, several limitations warrant attention.
Computational Cost. Although D3 is designed to be cost-aware, it remains substantially more expensive than single-judge evaluation. The D3-MORE protocol requires roughly four times the tokens of a single-judge setup, and D3-SAMRE can consume even more. This additional cost may be justified for high-stakes assessments or final validation but can be prohibitive for early-stage, iterative testing. The cost–accuracy frontier in Section 5.3 aims to make this trade-off explicit.
Persona Design and Fairness. The use of diverse juror personas, while empirically beneficial, risks reinforcing social or cultural stereotypes. Responsible persona construction demands careful auditing to ensure diversity does not compromise fairness or neutrality. Developing principled, automated methods for persona generation remains an open challenge in human–AI interaction and responsible AI research.
Dependence on Underlying Models. D3’s performance is ultimately constrained by the capabilities of the backbone LLM. Although its structure can elicit richer reasoning and reduce bias, it cannot introduce capabilities that the base model lacks. As LLMs improve, D3’s ceiling will rise correspondingly, but the dependency persists.
Scalability and Practical Use. Despite offering interpretability and robustness, D3’s multi-agent nature may limit its practicality for continuous evaluation pipelines. Future work could explore distilling D3’s rationale-rich judgments into smaller, specialized evaluator models or introducing game-theoretic interactions among agents to enhance efficiency without sacrificing rigor.
Appendix A Note on AI assistance
AI assistance was used for code development and improving the phrasing of the manuscript, while all analyses and conclusions were independently derived by the authors.
Appendix B Algorithms
Algorithm 1 Multi-Advocate One-Round Evaluation (MORE)
1:Initialize advocates $A_{1}=\{A_{11},...,A_{1k}\}$ for Answer 1, $A_{2}=\{A_{21},...,A_{2k}\}$ for Answer 2.
2:Initialize defenses $D_{1}\leftarrow\emptyset$, $D_{2}\leftarrow\emptyset$.
3:for $i=1$ to $k$ do $\triangleright$ Parallel argument generation
4:  $d_{1i}\leftarrow\text{GenerateArgument}(A_{1i},\text{Answer 1})$
5:  $D_{1}\leftarrow D_{1}\cup\{d_{1i}\}$
6:  $d_{2i}\leftarrow\text{GenerateArgument}(A_{2i},\text{Answer 2})$
7:  $D_{2}\leftarrow D_{2}\cup\{d_{2i}\}$
8:end for
[…]
11:Compile transcript $T$ with aggregated defenses.
[…]
13:for each Juror $C_{i}$ in panel do
14:  $v_{i}\leftarrow\text{Vote}(C_{i},T)$
15:  $V\leftarrow V\cup\{v_{i}\}$
16:end for
17:winner $\leftarrow\text{AggregateVotes}(V)$ $\triangleright$ Tie-break with Judge’s score if needed
18:return winner
Algorithm 2 Single Advocate Multi-Round Evaluation (SAMRE) with Budgeted Stopping
1:Initialize advocates $A_{1},A_{2}$, Judge $J$, Jurors $\{C_{1},...,C_{m}\}$, max rounds $R_{max}$, budget $B$.
2:Initialize transcript $T_{0}\leftarrow\emptyset$, scores $S\leftarrow[]$.
3:for $r=1$ to $R_{max}$ do
4:  $d_{1}^{r},d_{2}^{r}\leftarrow\text{GenerateArguments}(A_{1},A_{2},T_{r-1})$ $\triangleright$ Advocates update arguments
5:  $s_{1}^{r},s_{2}^{r},F^{r}\leftarrow\text{Evaluate}(J,d_{1}^{r},d_{2}^{r})$ $\triangleright$ Judge scores and gives feedback
6:  $S.\text{append}((s_{1}^{r},s_{2}^{r}))$
7:  $T_{r}\leftarrow T_{r-1}\cup\{d_{1}^{r},d_{2}^{r},s_{1}^{r},s_{2}^{r},F^{r}\}$
8:  if $\text{CheckConvergence}(S,\epsilon)$ or $\text{TokenCost}(T_{r})>B$ then
9:   break
10:  end if
11:end for
[…]
13:for $i=1$ to $m$ do
14:  $v_{i}\leftarrow\text{Vote}(C_{i},T_{r})$
15:  $V\leftarrow V\cup\{v_{i}\}$
16:end for
17:winner $\leftarrow\text{AggregateVotes}(V)$ $\triangleright$ Tie-break with Judge’s final score
18:return winner
Appendix C Proofs
C.1 Proof of Theorem 1 (Probabilistic Convergence)
Proof:
The theorem states that if $\lim_{r\to\infty}\mathbb{E}[\delta_{r}]=\Delta>0$, then $\delta_{r}$ converges in probability to $\Delta$. We want to show $\lim_{r\to\infty}P(|\delta_{r}-\Delta|<\epsilon)=1$ for any $\epsilon>0$.
We use the triangle inequality: $|\delta_{r}-\Delta|\leq|\delta_{r}-\mathbb{E}[\delta_{r}]|+|\mathbb{E}[\delta_{r}]-\Delta|$. For the event $\{|\delta_{r}-\Delta|\geq\epsilon\}$ to occur, it must be that either $\{|\delta_{r}-\mathbb{E}[\delta_{r}]|\geq\epsilon/2\}$ or $\{|\mathbb{E}[\delta_{r}]-\Delta|\geq\epsilon/2\}$.
By the assumption of convergence of the mean, for any $\epsilon>0$, there exists an $N_{1}$ such that for all $r\geq N_{1}$, $|\mathbb{E}[\delta_{r}]-\Delta|<\epsilon/2$. So the second condition does not hold for large $r$.
Now consider the first condition. By Chebyshev’s inequality:
$P(|\delta_{r}-\mathbb{E}[\delta_{r}]|\geq\epsilon/2)\leq\frac{\text{Var}(\delta_{r})}{(\epsilon/2)^{2}}=\frac{4\text{Var}(\delta_{r})}{\epsilon^{2}}.$
The variance of the Beta posterior is $\text{Var}(\delta_{r})=\frac{\alpha_{r}\beta_{r}}{(\alpha_{r}+\beta_{r})^{2}(\alpha_{r}+\beta_{r}+1)}$. Since $\alpha_{r}+\beta_{r}=\alpha_{0}+\beta_{0}+r$, the denominator grows as $O(r^{3})$, while the numerator $\alpha_{r}\beta_{r}$ grows at most as $O(r^{2})$. Thus, $\text{Var}(\delta_{r})=O(1/r)$, and $\lim_{r\to\infty}\text{Var}(\delta_{r})=0$.
Therefore, $\lim_{r\to\infty}P(|\delta_{r}-\mathbb{E}[\delta_{r}]|\geq\epsilon/2)=0$. Since both sources of deviation become arbitrarily small, $\lim_{r\to\infty}P(|\delta_{r}-\Delta|\geq\epsilon)=0$, which completes the proof. $\square$
C.2 Proof of Theorem 2 (Score-Separation via Parallel Advocacy)
Proof:
Let $g(f_{i,j})$ be the score of the $j$-th advocate for answer $a_{i}$. Let $G_{i}$ be the random variable representing the score of a single advocate for answer $a_{i}$. In the multi-advocate framework, the aggregated score is $M_{i}=\max(G_{i,1},...,G_{i,k})$.
We assume that answer $a_{1}$ is superior to $a_{2}$, formalized by stating that the cumulative distribution function (CDF) of $G_{1}$, denoted $F_{1}(x)$, first-order stochastically dominates (FOSD) the CDF of $G_{2}$, denoted $F_{2}(x)$. That is, $F_{1}(x)\leq F_{2}(x)$ for all $x$, and the inequality is strict for some $x$. This implies $\mathbb{E}[G_{1}]>\mathbb{E}[G_{2}]$.
The CDF of the maximum of $k$ i.i.d. samples from $G_{i}$ is $F_{M_{i}}(x)=(F_{i}(x))^{k}$. Since $F_{1}(x)\leq F_{2}(x)$ for all $x$, it follows that $(F_{1}(x))^{k}\leq(F_{2}(x))^{k}$. This means that $M_{1}$ also FOSD-dominates $M_{2}$, and thus $\mathbb{E}[M_{1}]>\mathbb{E}[M_{2}]$.
Furthermore, the operation of taking the maximum tends to stretch the upper tail of a distribution. The improvement from taking the maximum is expected to be greater for the stochastically larger distribution ($G_{1}$). Formally, $\mathbb{E}[M_{1}]-\mathbb{E}[G_{1}]\geq\mathbb{E}[M_{2}]-\mathbb{E}[G_{2}]$. This leads to a greater separation in expected scores:
$\mathbb{E}[M_{1}-M_{2}]=\mathbb{E}[M_{1}]-\mathbb{E}[M_{2}]>\mathbb{E}[G_{1}]-\mathbb{E}[G_{2}].$
This completes the proof. $\square$
C.3 Proof of Theorem 3 (Score Differentiation)
Proof:
Let $g(f_{i,j})$ be the score of the $j$-th advocate for answer $a_{i}$. Let $G_{i}$ be the random variable representing the score of a single advocate for answer $a_{i}$. In the multi-advocate framework, the aggregated score is $g(f_{i,agg})=\max_{j}g(f_{i,j})$. Let $M_{i}=\max(G_{i,1},...,G_{i,k})$ be the random variable for the aggregated score.
We assume that answer $a_{1}$ is superior to $a_{2}$. This can be formalized by stating that the cumulative distribution function (CDF) of $G_{1}$, denoted $F_{1}(x)$, is stochastically smaller than the CDF of $G_{2}$, denoted $F_{2}(x)$. That is, $F_{1}(x)\leq F_{2}(x)$ for all $x$, and there exists some $x$ for which the inequality is strict. This implies $\mathbb{E}[G_{1}]>\mathbb{E}[G_{2}]$.
The CDF of the maximum of $k$ i.i.d. samples from $G_{i}$ is $F_{M_{i}}(x)=(F_{i}(x))^{k}$. Since $F_{1}(x)\leq F_{2}(x)$, it follows that $(F_{1}(x))^{k}\leq(F_{2}(x))^{k}$. This means that $M_{1}$ is also stochastically larger than $M_{2}$, and thus $\mathbb{E}[M_{1}]>\mathbb{E}[M_{2}]$.
Furthermore, the operation of taking the maximum tends to stretch the upper tail of a distribution. The difference between the expected value of the maximum of $k$ samples and the expected value of a single sample is larger for distributions with more mass in the upper tail. Because $G_{1}$ is stochastically larger than $G_{2}$, the improvement from taking the maximum is expected to be greater for $a_{1}$.
$\mathbb{E}[M_{1}]-\mathbb{E}[G_{1}]\geq\mathbb{E}[M_{2}]-\mathbb{E}[G_{2}].$
This leads to a greater separation in expected scores:
$\mathbb{E}[M_{1}-M_{2}]=\mathbb{E}[M_{1}]-\mathbb{E}[M_{2}]>\mathbb{E}[G_{1}]-\mathbb{E}[G_{2}].$
This completes the proof. $\square$
Appendix D Notation and Scoring Criteria
D.1 Notation
  • $A=\{A_{1},A_{2}\}$: Set of advocates, where each advocate $A_{i}$ defends a specific answer.
  • $J$: The judge who evaluates the arguments presented by the advocates.
  • $C=\{C_{1},C_{2},C_{3}\}$: Set of jurors, where each juror $C_{i}$ casts a vote at the end of the evaluation process.
  • $s_{1}^{r}$ and $s_{2}^{r}$: Scores given by the judge in the $r$-th round, corresponding to the evaluations of $A_{1}$ and $A_{2}$, respectively.
  • $M_{r}$: The aggregated memory of all rounds up to the $r$-th round, which includes arguments, scores, and feedback.
  • $f_{A}(A,M_{r-1})$: Function that generates the arguments $a_{1}^{r}$ and $a_{2}^{r}$ for the advocates based on the previous memory $M_{r-1}$.
  • $f_{J}(J,a_{1}^{r},a_{2}^{r})$: Function that takes the judge and the arguments from the advocates, returning their scores $s_{1}^{r}$, $s_{2}^{r}$, and feedback $F^{r}$.
  • $f_{C_{i}}(C_{i},M_{r})$: Function that represents the voting decision made by each juror $C_{i}$ based on the final memory $M_{r}$.
  • $D_{i}$: The aggregated defense obtained by asking the LLM to consolidate the group’s defenses into a single summary.
D.2 Scoring Criteria
The judge scores the advocates’ arguments based on the following criteria, using a scale of 1-20:
  • Relevance to the question
  • Accuracy of information and use of credible sources
  • Depth of analysis and completeness of argument
  • Clarity of expression and logical flow
  • Strength of reasoning and factual support
  • Effectiveness in addressing opponent’s points
D.3 Juror Backgrounds
In the SAMRE design, we selected jurors with varied professional backgrounds and perspectives:
  • A retired professor of ethics
  • A young environmental activist
  • A middle-aged business owner
  • A social worker specializing in community development
  • A technology entrepreneur with a background in AI
Appendix E Data Preprocessing and Evaluation
E.1 Artifact Licensing and Availability
All benchmarks used in this study (MT-Bench, AlignBench, AUTO-J) are publicly available for research purposes under their respective licenses. Model APIs (GPT-4, Claude-3, Llama-3, Mistral) were accessed through their standard commercial or open-source terms of service. Baseline implementations follow the specifications in their original publications.
E.2 Data Preprocessing
To prepare the raw data for analysis, we implemented a script that processes the input data and generates an Excel file structured with the following columns:
  • Question: This column contains the aggregated user questions used for evaluation.
  • Response_A: This column includes the responses generated by Model A for each corresponding question.
  • Response_B: This column presents the responses generated by Model B for the same set of questions.
  • Model_A_Score: This binary score indicates the performance of Model A, where a score of 1 signifies a win and 0 signifies a loss in comparison to Model B.
  • Model_B_Score: Similarly, this binary score reflects the performance of Model B, with a score of 1 representing a win and 0 representing a loss against Model A.
This structured format allows for straightforward analysis and comparison of the models’ performances based on user questions and their respective responses.
Appendix F Agent Interaction Prompts
We provide the detailed prompts used for the interactions between agents in our proposed architecture. The prompts are designed to guide the agents effectively throughout the evaluation process.
F.1 Multi-Advocate One-Round Evaluation (MORE) Architecture Prompts
F.1.1 Judge Prompt
You’re a critical, impartial judge in a high-stakes debate on: "{question}".
Answer 1: "{answer1}". Answer 2: "{answer2}".
Your goal is to provide detailed, constructive feedback that will push advocates to significantly improve their arguments.
Current round: {current_round}
Max rounds: {max_rounds}
Previous scores: {previous_scores}
Defense for 1st answer: {defense1}
Defense for 2nd answer: {defense2}
Analyze each argument meticulously. Be thorough and unbiased in your assessment of:
1. Relevance to the question
2. Accuracy of information and use of credible sources
3. Depth of analysis and completeness of argument
4. Clarity of expression and logical flow
5. Strength of reasoning and factual support
6. Effectiveness in addressing opponent’s points
For each criterion, provide a score on a scale of 1-20 and detailed justification.
Scores should be given as [Answer1_score, Answer2_score] for each criterion.
Your comprehensive feedback for each advocate (50 words each):
Feedback for Advocate 1:
Feedback for Advocate 2:
Sum up the scores and return the final score tuple (score1, score2). Example: (95, 87)
Your detailed scores and final tally:
F.1.2 Advocate Prompts
You’re a fierce advocate defending this answer: {answer} to the question: {question}. Your opponent defends: {opponent_answer}. Convince the audience your argument is superior by addressing these criteria:
1. Relevance
2. Accuracy
3. Depth
4. Clarity
5. Logic and Factuality
Be assertive and don’t hold back! Use rhetoric and persuasion to win over the audience but be respectful.
Latest feedback: {feedback}
Opponent’s last argument: {opponent_argument}
Your fierce defense (100 words max):
F.1.3 Summarizer Prompt
[…]
F.2 Single Advocate Multi-Round Evaluation (SAMRE)
F.2.1 Defend Answer Prompt
You’re an eloquent advocate (Advocate {advocate_id}) in a group defending this answer: {answer} to the question: {question}. Your opponent group defends: {opponent_answer}. Collaborate with your fellow advocates to convince the audience your argument is better. Use the latest feedback, your opponent’s last argument, and your team’s previous arguments to improve your case.
Latest feedback: {feedback}
Opponent’s last argument: {opponent_argument}
Your team’s previous arguments: {team_arguments}
Respond in under 80 words.
Your defense:
F.2.2 Aggregate Defense Prompt
You are an expert debate strategist. Your task is to aggregate and improve upon the following defenses for the answer: {answer} to the question: {question}. The opponent’s answer is: {opponent_answer}.
Individual defenses: {defenses}
Latest feedback from the judge: {feedback}
Analyze each defense critically. Identify the strongest points, address any weaknesses, and combine the best arguments into a cohesive, powerful defense. Aim to create a defense that is stronger and more comprehensive than any individual argument.
Provide your aggregated and improved defense in under 150 words:
F.2.3 Judge Answer Prompt
You’re a fair, impartial judge in a debate on: "{question}". Answer 1: "{answer1}". Answer 2: "{answer2}". Your goal is to provide feedback that will help advocate groups improve and differentiate their arguments more clearly.
Current round: {current_round}
Total rounds: {total_rounds}
Previous scores: {previous_scores}
Defense for 1st answer: {defense1}
Defense for 2nd answer: {defense2}
Provide specific, constructive feedback to help each advocate group strengthen their unique position. Encourage them to address weaknesses and highlight distinctions. Aim for your feedback to lead to more divergent scores in future rounds.
Give your feedback in under 50 words:
F.2.4 Score Answer Prompt
You’re a critical, impartial judge in a high-stakes debate on: "{question}". Answer 1: "{answer1}". Answer 2: "{answer2}". Your goal is to provide detailed, constructive feedback that will push advocates to significantly improve their arguments.
Total rounds: {total_rounds}
Previous scores: {previous_scores}
Defense for 1st answer: {defense1}
Defense for 2nd answer: {defense2}
Analyze each argument meticulously. Be thorough and unbiased in your assessment of:
1. Relevance to the question
2. Accuracy of information and use of credible sources
3. Depth of analysis and completeness of argument
4. Clarity of expression and logical flow
5. Strength of reasoning and factual support
6. Effectiveness in addressing opponent’s points
For each criterion, provide a score on a scale of 1-20 and detailed justification. Scores should be given as [Answer1_score, Answer2_score] for each criterion.
Your comprehensive feedback for each advocate (50 words each):
Feedback for Advocate 1:
Feedback for Advocate 2:
Sum up the scores and return the final score tuple (score1, score2). Example: (95, 87)
Your detailed scores and final tally:
F.3 Baseline Model Prompt
You are a fair, impartial judge scoring a debate on the following question: {question}.
Answer 1: {answer1}
Answer 2: {answer2}
Score each answer on a scale of 1-20 for each of the following criteria:
1. Relevance to the question
2. Accuracy of information and use of credible sources
3. Depth of analysis and completeness of argument
4. Clarity of expression and logical flow
5. Strength of reasoning and factual support
6. Effectiveness in addressing opponent’s points
Provide scores as [Answer1_score, Answer2_score] for each criterion in a list format, then sum for final scores. Please keep an eye on the slightest difference that should make a difference in the scoring. Don’t overthink!
Relevance:
Accuracy:
Depth:
Clarity:
Logic and Factuality:
Addressing opponent’s points:
Final Scores (sum of above) as a tuple (example: (18, 9)):
Explain your scoring, focusing on why one answer is better than the other based on the criteria above. Keep your explanation concise but informative.
Finally, return the final score tuple (score1, score2) as a tuple (in parentheses). Example: (18, 9)
Your scores and explanation:

### v4

@@
 (2023), AlignBench Liu et al.
-(2024), and AUTO-J, showing state-of-the-art agreement with human judgments (accuracy and Cohen’s $\kappa$), reduced positional and verbosity biases via anonymization and role diversification, and a favorable cost–accuracy frontier enabled by budgeted stopping.
+(2024), and AUTO-J Li et al.
+(2023), showing state-of-the-art agreement with human judgments (accuracy and Cohen’s $\kappa$), reduced positional and verbosity biases via anonymization and role diversification, and a favorable cost–accuracy frontier enabled by budgeted stopping.
 Ablations and qualitative analyses isolate the contributions of debate, aggregation, and anonymity.
-Together, these results establish D3 as a principled, practical recipe for reliable, interpretable, and cost-aware LLM evaluation.
+Together, these results establish D3  [^1] Code Available at: https://github.com/abirharrasse/D3-Judge as a principled, practical recipe for reliable, interpretable, and cost-aware LLM evaluation.
 Debate, Deliberate, Decide (D3): A Cost-Aware Adversarial Framework for Reliable and Interpretable LLM Evaluation
-Chaithanya Bandi1   Abir Harrasse1,2 1Martian  2Mohammed VI Polytechnic University {chaithanya@withmartian.com, abir.harrasse@emines.um6p.ma}
+Abir Harrasse1,   Chaithanya Bandi1,2,   Hari Bandi3 1Martian  2NUS  3MIT {abir.harrasse@emines.um6p.ma, bizchaba@nus.edu.sg}
+Figure 1: The D3 system routes pairwise evaluations through structured adversarial debate.
+Two protocols (MORE for efficiency and SAMRE with budgeted stopping for depth) generate parallel or iterative arguments.
+A Judge provides criteria-based scoring to guide refinement.
+A diverse jury panel independently evaluates the anonymized debate transcript and renders a final verdict, with ties broken by the Judge’s cumulative score.
 1 Introduction
 The rapid proliferation of Large Language Models (LLMs) Brown et al.
-(2020) has created a significant challenge in evaluating their increasingly complex and nuanced capabilities, particularly in open-ended generation tasks Celikyilmaz et al.
+(2020) has created significant challenges in evaluating their increasingly complex capabilities, particularly in open-ended generation tasks Celikyilmaz et al.
 (2021).
 Traditional automated metrics such as BLEU Papineni et al.
-(2002) and ROUGE Lin (2004), while useful for specific tasks like translation and summarization, often fail to capture semantic coherence, factual accuracy, or alignment with human values Callison-Burch et al.
+(2002) and ROUGE Lin (2004) fail to capture semantic coherence, factual accuracy, or alignment with human values Callison-Burch et al.
 (2006); Mathur et al.
 (2020).
-Consequently, human evaluation has remained the gold standard Howcroft et al.
-(2020), but its high cost, slow turnaround, and inherent subjectivity make it impractical for the iterative development cycles of modern AI systems Liang et al.
+Consequently, human evaluation remains the gold standard Howcroft et al.
+(2020), but its high cost, slow turnaround, and inherent subjectivity make it impractical for iterative development cycles Liang et al.
 (2023).
-This scalability crisis has given rise to the “LLM-as-a-Judge” paradigm, where a powerful LLM is used to evaluate the outputs of other models, showing promising alignment with human preferences Zheng et al.
+This has motivated the “LLM-as-a-Judge” paradigm Zheng et al.
 (2023); Kim et al.
-(2024).
-Such evaluation approaches are critical for training helpful and harmless assistants through reinforcement learning from human feedback Bai et al.
+(2024), where powerful LLMs evaluate other models’ outputs, showing promising alignment with human preferences.
+Such approaches are critical for training helpful assistants via reinforcement learning from human feedback Bai et al.
 (2022); Christiano et al.
 (2023); Ziegler et al.
-(2020), building general language assistants aligned with human values Askell et al.
-(2021), and benchmarking comprehension capabilities Rajpurkar et al.
+(2020) and building aligned language assistants Askell et al.
 (2021).
-However, the reliance on a single LLM judge introduces its own set of critical vulnerabilities.
-These systems are susceptible to a range of biases, including positional bias (favoring the first or second response presented), verbosity bias (preferring longer answers), and self-enhancement bias (unfairly favoring outputs from its own model family) Wang et al.
+However, single LLM judges are susceptible to positional bias, verbosity bias, and self-enhancement bias Wang et al.
 (2023); Mehrabi et al.
 (2022).
-To address these shortcomings, the research community has turned to multi-agent systems, which leverage the collective intelligence of multiple LLM agents to produce more robust and reliable evaluations.
-Foundational frameworks like ChatEval Chan et al.
-(2024) have demonstrated that orchestrating debates among agents with diverse personas can improve correlation with human judgments Li et al.
-(2024); Hong et al.
-(2024).
-Similarly, PRD (Peer Rank & Discussion) Li et al.
-(2024) introduced mechanisms for peer review and discussion to mitigate biases and reach a more balanced consensus Wang et al.
-(2024).
-Despite the promise of these multi-agent approaches, their widespread adoption has been hindered by several persistent gaps in the existing research Celikyilmaz et al.
-(2021).
-First, many studies suffer from insufficient empirical rigor, often confining their validation to a single benchmark and failing to compare against a comprehensive suite of strong, contemporary baselines Liang et al.
-(2023).
-Second, while these frameworks often claim to reduce bias, they frequently lack a dedicated methodology for auditing and quantifying specific biases, leaving their robustness as an unverified assertion Wang et al.
-(2023).
-Third, and perhaps most critically for practical application, most frameworks are cost-agnostic in their design.
-The computational cost is a significant barrier to adoption, yet it is rarely treated as a primary design constraint or a formal evaluation metric Zheng et al.
-(2023).
-This paper introduces the Debate, Deliberate, Decide (D3) framework, a next-generation multi-agent evaluation system designed explicitly to address these critical gaps.
-D3 is built upon a foundation of empirical rigor and practical applicability, offering a principled approach to LLM evaluation that is both reliable and cost-aware.
-The contributions of this work are fourfold:
+Multi-agent approaches like ChatEval Chan et al.
+(2024) and PRD Li et al.
+(2024) mitigate these through diverse personas and peer discussion, yet critical gaps remain: insufficient empirical rigor across diverse benchmarks, lack of dedicated bias auditing methodologies, and cost-agnostic designs despite computational expense being a fundamental adoption barrier.
+This paper introduces Debate, Deliberate, Decide (D3), a multi-agent evaluation system addressing these gaps through a uniquely integrated approach.
+Our contributions are:
   1. 1.
-    A modular, multi-agent architecture (D3) featuring specialized Advocate, Judge, and Juror roles with two distinct, cost-aware debate protocols: a parallelized Multi-Advocate One-Round (MORE) protocol for efficiency and an iterative Single-Advocate Multi-Round (SAMRE) protocol for depth.
+    Courtroom-inspired architecture with explicit Advocate/Judge/Juror role specialization and structured adversarial debate protocols.
   2. 2.
-    A theoretically grounded reliability model that formalizes the error-reduction dynamics of iterative debate, justifying a budgeted stopping rule that optimizes the trade-off between evaluation confidence and computational cost.
+    Dual cost-aware protocols: MORE for parallelized efficiency and SAMRE with budgeted stopping for iterative depth, offering the first explicit cost-accuracy trade-off menu for practitioners.
   3. 3.
-    A comprehensive empirical validation across three diverse benchmarks: MT-Bench Zheng et al.
-(2023) for conversational ability, AlignBench Liu et al.
-(2024) for alignment, and the AUTO-J dataset for real-world scenario coverage, against a suite of strong baselines including ChatEval Chan et al.
-(2024), PRD Li et al.
-(2024), and the fine-tuned evaluator PandaLM Wang et al.
-(2024).
+    Theoretical grounding via probabilistic convergence models and formal separation proofs that justify design choices and stopping criteria.
   4. 4.
-    A systematic bias auditing methodology that quantifies and demonstrates D3’s superior robustness to positional and self-enhancement biases through controlled experiments Wang et al.
-(2023); Mehrabi et al.
-(2022).
-By systematically addressing the limitations of prior work, D3 provides a scalable, interpretable, and cost-sensitive solution that advances the state of the art in automated LLM evaluation, paving the way for more trustworthy and reliable AI systems.
+    Systematic bias auditing methodology that quantifies and demonstrates robustness to positional and self-enhancement biases through controlled experiments.
+  5. 5.
+    Rigorous multi-benchmark validation across MT-Bench, AlignBench, and AUTO-J against strong baselines (ChatEval, PRD, PandaLM).
+By combining all five dimensions, D3 provides a scalable, interpretable, and cost-sensitive solution that advances the state of the art in trustworthy LLM evaluation.
 2 The Debate, Deliberate, Decide (D3) Framework
@@
 The hypothesis is that persona diversity allows the evaluation to capture a wider range of qualitative aspects, leading to a decision that is better aligned with a broad spectrum of human values.
+We validated the robustness of persona selection by testing 50 diverse personas across varying domains; details in Appendix G.1.
 2.2 The Adversarial Debate Protocols
@@
 4.1 Benchmarks and Evaluation Tasks
-To assess generalizability, we evaluate on three benchmarks targeting different LLM capabilities: MT-Bench (80 multi-turn questions), AlignBench (683 questions translated to English), and AUTO-J (58 scenarios, 3,436 pairwise comparisons).
-  • MT-Bench: A standard benchmark for evaluating multi-turn conversational and instruction-following capabilities across a range of common user queries [1, 1].
-It serves as a test of general-purpose helpfulness.
-  • AlignBench: A benchmark specifically designed to evaluate the alignment of LLMs with human values, covering dimensions such as helpfulness, harmlessness, and ethical considerations in Chinese contexts.[9, 10] For this study, the queries were professionally translated to English to test alignment capabilities beyond a single language.
-  • AUTO-J Dataset: A comprehensive dataset featuring 58 real-world scenarios, with pairwise preferences judged by GPT-4.[11, 12] This benchmark tests the evaluator’s ability to generalize across a wide variety of user intents and task domains, from creative writing to technical explanation, providing a strong test of real-world applicability.
+We evaluate on three benchmarks targeting different LLM capabilities:
+  • MT-Bench: 80 multi-turn conversational questions testing general-purpose helpfulness and instruction-following Zheng et al. (2023).
+  • AlignBench: 683 alignment-focused questions covering helpfulness, harmlessness, and ethical reasoning (professionally translated to English) Liu et al. (2024).
+  • AUTO-J: 58 real-world scenarios with 3,436 pairwise comparisons spanning creative writing, technical explanation, and diverse task domains Li et al. (2023).
+Figure 2: Persona Effectiveness by Task Type. Personas improve performance on subjective tasks (+5.7%) but provide minimal gains on objective tasks (+0.5%).
 4.2 Models and Comparative Baselines
-The experiments involved a diverse set of LLMs, both as generators of content to be evaluated and as the agents within the evaluation frameworks themselves.
-  • Models Under Evaluation: To generate the candidate answers for evaluation, a set of leading proprietary and open-source models were used, including GPT-4-Turbo, Claude-3-Opus, Llama-3-70B, and Mistral-Large.
-All model versions were explicitly logged to ensure reproducibility.
-  • Evaluator Models: To provide a strong and consistent foundation for all evaluation frameworks, GPT-4-Turbo was used as the backbone model for all agent roles (Advocate, Judge, Juror) in D3 and in the baseline implementations.
-To assess the framework’s adaptability and provide a more cost-effective option, a parallel set of experiments was run using Llama-3-70B as the evaluator model.
-  • Baselines: The D3 framework was compared against a carefully selected set of four strong baselines to provide a comprehensive picture of its relative performance.
-    1. 1.
-      Single LLM-as-a-Judge: The standard baseline where a single instance of GPT-4-Turbo is prompted to directly choose the better of two answers, reflecting the most common approach in practice.
-    2. 2.
-      ChatEval: A leading multi-agent debate framework that uses agents with diverse personas to discuss and score responses.
-The implementation followed the official description of the method to ensure a fair comparison.[5, 6]
-    3. 3.
-      PRD (Peer Rank & Discussion): Another influential multi-agent baseline that leverages peer discussion to reach a consensus, specifically designed to mitigate biases like self-enhancement.[4, 8]
-    4. 4.
-      PandaLM: A powerful open-source evaluator model that has been specifically fine-tuned for the task of judging LLM outputs.[13, 14] This baseline represents the state-of-the-art in specialized evaluator models and provides a comparison point against a non-debate-based approach.
+We employ a diverse set of LLMs across two roles.
+For content generation, we use GPT-4-Turbo OpenAI (2024), Claude-3-Opus Anthropic (2025), Llama-3-70B Llama (2024), and Mistral-Large Mistral AI (2024).
+For evaluation, GPT-4-Turbo serves as the backbone LLM for all agent roles (Advocate, Judge, Juror) in D3 and baseline implementations, ensuring fair comparison.
+We additionally run experiments with Llama-3-70B Llama (2024) as evaluator to assess framework performance with open-source models and provide cost-effective alternatives.
+We compare D3 against four strong baselines: (1) a single GPT-4-Turbo judge directly selecting the better answer, representing standard practice; (2) ChatEval, a leading multi-agent framework using diverse personas to debate and score responses; (3) PRD (Peer Rank & Discussion), which leverages peer-review mechanisms to mitigate self-enhancement and positional biases; and (4) PandaLM, a specialized fine-tuned evaluator representing state-of-the-art in non-debate approaches.
 4.3 Core Metrics and Bias Audits
-A multi-faceted evaluation was conducted using metrics that capture not only accuracy but also efficiency and fairness.
-  • Agreement with Human Judgments:
-    • Accuracy: The primary metric, defined as the percentage of times the evaluator framework’s final verdict matches the ground-truth human preference provided by the benchmark.
-    • Cohen’s Kappa ($\kappa$): This metric measures inter-rater agreement while correcting for the probability of agreement occurring by chance.
-It provides a more robust measure of consensus than raw accuracy, especially on datasets with skewed preference distributions.
-  • Efficiency:
-    • Average Tokens per Evaluation: The total number of input and output tokens consumed by all API calls required to produce a single pairwise judgment, averaged across the entire test set.
-This serves as a direct proxy for computational cost and operational expense.
-  • Bias Auditing:
-    • Positional Swap Consistency: To measure positional bias, each evaluation was performed twice: once with the answers in the order (A, B) and once with the order (B, A).
-Consistency is the percentage of evaluations where the final verdict was invariant to this swap (e.g., if A was preferred in the first run, B was preferred in the second).
-A higher score indicates lower positional bias.[11]
-    • Self-Enhancement Rate: This audit was conducted on a subset of evaluations where one of the candidate answers was generated by a model from the same family as the evaluator agent (e.g., GPT-4-Turbo judging a response from another GPT-4 variant).
-The rate is the percentage of instances where the evaluator’s preference for its own family’s output disagreed with the human label, indicating potential favoritism.[3]
+We measure accuracy and Cohen’s Kappa ($\kappa$) for agreement with human judgments, with Kappa correcting for chance agreement on skewed distributions.
+Efficiency is measured as average tokens per evaluation (proxy for computational cost).
+Bias audits include: (1) Positional Swap Consistency: each evaluation performed twice with answers in order (A, B) and (B, A), measuring consistency of verdicts; (2) Self-Enhancement Rate: percentage of cases where evaluator prefers its own model family despite human labels indicating otherwise, measured on subset where one answer is from same model family as evaluator.
 5 Results and In-Depth Analysis
+5.1 Persona Sensitivity Analysis
+We validate our persona design through comprehensive ablation studies.
+We created a diverse pool of 50 personas spanning law, medicine, education, technology, ethics, business, social work, risk analysis, and compliance.
+We conducted 10 experiments using random 5-persona subsets on MT-Bench (120 questions), observing $85.9\%\pm 0.7\%$ accuracy, compared to $86.1\%$ with our curated set and $82.7\%\pm 1.1\%$ with generic jurors.
+This demonstrates that persona conditioning provides a consistent $3.2\%$ gain ($p<0.01$) while remaining robust to specific persona choice.
+Our curated set was designed to provide complementary expertise and value perspectives: ethics and human values (ethics professor); social and environmental impact (environmental activist, social worker); business and practical trade-offs (business owner); and technology and innovation (tech entrepreneur).
+Personas influence attention and reasoning style rather than demographics, avoiding stereotyping, with all jurors sharing the same backbone LLM.
+5.2 Task-Type Analysis: Objective vs. Subjective Tasks
+We performed category-level analysis on MT-Bench to understand when personas provide value.
+On objective tasks (coding, math, factual QA), personas provide minimal benefit: $80.6\%$ vs.
+$80.1\%$ for D3-MORE without personas ($0.5\%$).
+On subjective tasks (writing, reasoning, ethics, roleplay), personas deliver substantial gains: $80.9\%$ vs.
+$75.2\%$ ($5.7\%$).
+This validates that D3 excels where diverse perspectives and value alignment matter most, with personas being most effective on value-laden evaluations (Figure 2).
 5.3 D3 Achieves State-of-the-Art Agreement with Human Judgments
-As shown in Table 1, both variants of the D3 framework consistently and significantly outperform all baselines across the three diverse benchmarks.
-The D3-MORE protocol, designed for efficiency, surpasses the next best baseline, ChatEval, by a significant margin on all datasets.
-For instance, on MT-Bench, D3-MORE achieves an accuracy of 85.1%, representing a 12.6% absolute improvement over the standard Single Judge baseline and a 6.9% improvement over ChatEval.
+As shown in Table 1, both variants of the D3 outperform all baselines across the three diverse benchmarks.
+The D3-MORE protocol, designed for efficiency, surpasses the next best baseline, ChatEval Chan et al.
+(2024), by a significant margin on all datasets.
+For instance, on MT-Bench Zheng et al.
+(2023), D3-MORE achieves an accuracy of 85.1%, representing a 12.6% absolute improvement over the standard Single Judge baseline and a 6.9% improvement over ChatEval.
 The D3-SAMRE protocol, which allows for deeper iterative refinement, achieves the highest overall accuracy, reaching 86.3% on MT-Bench.
@@
 Table 1: Main performance comparison of evaluation frameworks. D3 variants demonstrate superior agreement with human judgments across all three benchmarks in both accuracy and Cohen’s Kappa.
-5.2 Auditing and Mitigating Evaluation Biases
-A core design goal of D3 was to create a system that is not only accurate but also fair and robust.
-The bias audit results, presented in Table 2, confirm the success of this objective.
-The Single Judge baseline exhibits a notable susceptibility to positional bias, with a Positional Swap Consistency of only 81.7%, meaning its judgment changes in nearly one out of five cases simply by swapping the order of the answers.
-While other multi-agent baselines like ChatEval and PRD offer improvements, the D3 framework demonstrates exceptional robustness, with D3-SAMRE achieving a consistency of 96.2%.
-This resilience is attributable to the structured debate and the multi-juror deliberation process, which forces a judgment based on the merit of the arguments rather than superficial cues.
-Furthermore, D3 shows a markedly lower Self-Enhancement Rate (8.4%) compared to the Single Judge (24.6%), indicating that the anonymized, adversarial nature of the debate effectively mitigates the tendency of an LLM to favor its own outputs.
-Framework Positional Swap Consist. (%) Self-Enhancement Rate (%)
-Single Judge 81.7 24.6
-ChatEval 88.5 14.1
-PRD 89.2 12.5
-PandaLM 92.4 N/A*
-D3-MORE (Ours) 94.8 9.1
-D3-SAMRE (Ours) 96.2 8.4
-Table 2: Bias audit results.
-D3 exhibits significantly higher consistency when answer positions are swapped and a lower rate of unfairly favoring its own model family’s outputs compared to baselines.
-(*PandaLM is a Llama-based model; self-enhancement was not measured against the proprietary models used for generation in this test.)
+We further evaluate D3 using a fully open-source model Llama3-70B Llama (2024) to verify that its gains are not specific to proprietary models; detailed results are reported in Appendix H.
 5.4 Characterizing the Cost-Versus-Accuracy Frontier
-A critical contribution of this work is the formal analysis of the trade-off between evaluation cost and accuracy.
-By plotting the accuracy of each framework against its average token consumption, a clear cost-accuracy frontier emerges.
-The D3 framework establishes a new, superior Pareto frontier, offering better performance at multiple cost points.
-For example, D3-MORE achieves substantially higher accuracy than all baselines, including ChatEval and PRD, while maintaining a competitive token cost.
-This makes it an ideal choice for large-scale evaluations where budget is a primary concern.
-D3-SAMRE, with its budgeted stopping rule, provides a high-end option that delivers the highest possible accuracy for users who can afford a larger computational budget.
-This analysis moves the field beyond a singular focus on accuracy, providing practitioners with a principled and data-driven way to select an evaluation method that aligns with both their quality requirements and their operational constraints.
-5.4 Ablation Studies: Deconstructing the Framework’s Efficacy
-To empirically validate the specific design choices within the D3 architecture, a series of ablation studies were conducted on the MT-Bench dataset.
-The results, summarized in Table 3, demonstrate that each component contributes meaningfully to the framework’s overall performance.
-  • The Value of Juror Diversity: When the diverse juror personas were replaced with a homogeneous panel of generic "impartial evaluators," the accuracy of D3-MORE dropped by 3.8 percentage points, and its positional consistency decreased.
-This result provides strong evidence that persona diversity is not a superficial feature but a critical mechanism for reducing correlated errors and achieving a more robust consensus.
-  • The Impact of Iterative Refinement: Comparing the full D3-SAMRE protocol to a version that was artificially limited to a single round of debate revealed a 2.1% drop in accuracy.
-This confirms that the iterative process is crucial for resolving difficult cases between closely matched answers, justifying its higher token cost.
-  • The Role of the Judge: A variant of D3 was tested where advocates debated without the structured feedback and scoring of a Judge, with the transcript sent directly to the jurors.
-This led to a significant degradation in performance, as the debate lacked focus and failed to systematically address key criteria.
-This highlights the Judge’s essential role in moderating the debate and ensuring a high-quality, evidence-based discussion.
-D3 Variant Acc. (%) on MT-Bench Pos. Swap Consist. (%) Avg. Tokens
-Full D3-MORE 85.1 94.8 $\sim$5,000
-D3-MORE w/o Diverse Jurors 81.3 90.1 $\sim$5,000
-Full D3-SAMRE 86.3 96.2 $\sim$7,000
-D3-SAMRE (1 Round Only) 84.2 93.5 $\sim$4,800
-D3 (MORE) w/o Judge 79.5 88.7 $\sim$4,500
-Table 3: Ablation study of D3 components.
-The results show that juror diversity, iterative refinement, and the presence of a moderating Judge all provide statistically significant contributions to the framework’s performance and robustness.
-5.5 Qualitative Analysis
-To illustrate the practical difference in evaluation quality, consider a query from AlignBench asking for a balanced discussion on the ethical implications of AI in hiring.
-The Single Judge preferred a verbose but generic response that listed potential pros and cons without depth.
-In contrast, the D3 framework correctly identified a more nuanced answer as superior.
-During the debate, the advocate for the better answer argued: "Response B moves beyond a simple list to discuss the disparate impact on protected groups and suggests concrete mitigation strategies, demonstrating a deeper understanding of fairness." One juror, with the persona of an ethics professor, wrote in their rationale: "Response A is descriptive, but Response B is prescriptive and aligned with principles of justice.
-It doesn’t just identify problems; it engages with solutions.
-The choice is clear." This example showcases how the D3 process—through adversarial argumentation and diverse expert perspectives—can uncover qualitative distinctions that a single, monolithic evaluation might miss.
+D3-MORE achieves $85.1\%$ accuracy at $\sim 3,050$ tokens ($\mathdollar 0.31$), delivering $6.9\%$ higher accuracy than ChatEval at comparable cost (35.8 tokens/accuracy point) (Figure 3).
+It also outperforms PRD by $8.3\%$ accuracy at lower cost.
+D3-SAMRE reaches $86.3\%$ accuracy with mean cost of $4.65\times$ single judge due to iterative debate; however, 58% of debates stop by round 2 via budgeted stopping, with actual mean of 2.71 rounds (not maximum 5).
+MORE is highly parallelizable with latency comparable to ChatEval; SAMRE is sequential but suitable for batch evaluation.
+This analysis provides practitioners a principled way to select between efficiency (D3-MORE) and peak accuracy (D3-SAMRE) aligned with quality and budget constraints.
+See Table 2 for complete comparison.
+Figure 3: The Cost-Accuracy Pareto Frontier.
+D3-MORE and D3-SAMRE (with budgeted stopping) establish a new frontier, providing higher human agreement at lower or comparable costs than existing multi-agent baselines like ChatEval.
+Framework Accuracy (%) Avg. Tokens Tokens/Acc. Cost
+Single Judge 72.5 850 11.7 \$0.09
+PandaLM 75.5 1,100 14.6 \$0.11
+ChatEval 78.2 2,800 35.8 \$0.28
+PRD 76.8 3,100 40.4 \$0.31
+D3-MORE 85.1 3,050 35.8 \$0.31
+D3-SAMRE 86.3 4,650 53.9 \$0.47
+Table 2: Cost-Accuracy Analysis: D3 Framework Performance vs. Baselines (MT-Bench)
+5.5 Protocol Selection Guidelines
+We provide empirical guidance for selecting between D3 protocols.
+SAMRE improves over MORE in 28% of cases but incurs +42% token cost; MORE remains correct when SAMRE fails in only 4% of cases.
+SAMRE excels on multi-turn reasoning, ethical trade-offs, roleplay requiring nuance, and cases with initial judge score gaps <25<25 points.
+We suggest to use SAMRE for high-stakes evaluations, borderline quality cases, or ethically complex scenarios requiring deeper refinement and to use MORE as the efficient default for all other cases.
+5.6 Disentangling Ensemble Effects from Persona-Based Gains
+We conduct a 4-way ablation study on MT-Bench to decompose the contributions of multi-juror ensembles and persona-based evaluation (Table 3).
+Configuration Accuracy (%) Pos. Swap Consist. (%) Cohen’s $\kappa$
+Single juror, no persona 72.5 81.7 0.45
+Single juror, with persona 74.8 83.2 0.47
+Multi-juror (k=5), no personas 81.3 90.1 0.54
+Multi-juror (k=5), with personas 85.1 94.8 0.58
+Table 3: Ablation Study. Decomposing Ensemble and Persona Effects on MT-Bench
+  • Ensemble Effect Dominates: Scaling from a single juror to five jurors yields the largest gain (+8.8%), demonstrating that diversity of judgment is the primary driver of D3-MORE’s performance.
+Multi-juror consensus reduces correlated errors and improves both accuracy and consistency metrics (positional swap consistency: 81.7% → 90.1%, Cohen’s $\kappa$: 0.45 → 0.54).
+  • Personas Provide Consistent Multiplicative Gains: Beyond ensemble effects, persona-based evaluation adds +3.8% accuracy (statistically significant, $p<0.01$).
+Critically, this persona effect is consistent across all jury sizes (k=1, 3, 5, 7), with personas providing a stable 3–4% boost independent of ensemble size.
+  • Synergistic Interaction: The combination of ensembles and personas (85.1%) exceeds their individual contributions, with personas enabling smaller juries to match larger homogeneous ones (k=5 with personas $\approx$ k=7 without personas, $85.1\%$ vs.
+$83.2\%$), suggesting efficient allocation of computational resources.
+All comparisons include 95% confidence intervals via bootstrap resampling ($n=1000$) and paired $t$-tests for statistical significance.
+5.7 Budgeted Stopping and Cost Efficiency Analysis
+Analysis of 1,200 SAMRE evaluations shows 58% converge by round 2 (Figure 4), maintaining 92% accuracy while reducing token consumption 40% versus fixed 5-round debates.
+Forced continuation beyond convergence changes verdicts in only 6% of cases, primarily tie scenarios.
+This demonstrates D3 effectively identifies diminishing returns in evaluation depth, validating the budgeted stopping rule as a practical mechanism for cost control without sacrificing reliability.
+Figure 4: SAMRE evaluation demonstrates that budgeted stopping reduces token consumption by 40% without compromising the 92% accuracy achieved by fixed protocols, validating the approach as a practical mechanism for cost control.
+5.8 Interpretability Analysis
+We quantify interpretability through concrete metrics on 500 juror rationales.
+  1. 1.
+    Evidence Citation Rate: 94% of rationales explicitly reference debate transcripts (e.g., “Advocate A argued that…”, “In round 3, the rebuttal addressed…”).
+  2. 2.
+[…]
+  3. 3.
+    Disagreement Traceability: among 50 sampled juror disagreements, 96% had clearly identifiable reasoning axes (e.g., accuracy vs. comprehensiveness, safety vs. helpfulness).
+We define interpretability in a process-centric sense: D3 provides structured debate transcripts with attributed arguments and juror rationales, making evaluation decisions transparent and auditable.
+This contrasts with black-box judges and enables practitioners to verify verdicts align with stated evaluation criteria.
 6 Related Work
 6.1 LLM-as-a-Judge
-The concept of using a powerful LLM to evaluate the outputs of other models has become a cornerstone of modern LLM assessment Brown et al.
+Using a strong language model to evaluate the outputs of other models has become a central paradigm in modern LLM assessment Brown et al.
 (2020).
-This paradigm was popularized by studies demonstrating that models like GPT-4 can achieve high agreement with human preferences, offering a scalable alternative to manual annotation Zheng et al.
+Prior work has shown that models such as GPT-4 can achieve high agreement with human preferences, enabling scalable evaluation without extensive manual annotation Zheng et al.
 (2023); Kim et al.
 (2024).
-Benchmarks such as MT-Bench and Chatbot Arena are built upon this principle, collecting pairwise comparisons to establish leaderboards of model performance Zheng et al.
+Widely used benchmarks including MT-Bench and Chatbot Arena operationalize this paradigm through pairwise comparisons judged by a single LLM Zheng et al.
 (2023).
-While foundational, this single-judge approach is the primary target for the improvements D3 aims to deliver, specifically addressing its documented vulnerabilities to cognitive and positional biases Wang et al.
+Despite its practicality, the single-judge paradigm is known to suffer from systematic failure modes, including positional bias, verbosity bias, and preference leakage from prompt structure Wang et al.
 (2023); Mehrabi et al.
 (2022).
+Recent work has further shown that LLM judges may inadvertently encode information about the evaluated model or task context, leading to biased or unstable evaluations Li et al.
+(2025).
+D3 targets this setting directly, treating single-judge evaluation as a strong but brittle baseline and explicitly addressing its bias and robustness limitations through structured deliberation and anonymization.
 6.2 Multi-Agent Debate for Evaluation
-To overcome the limitations of a single judge, researchers have explored multi-agent systems that simulate collaborative human evaluation processes Hong et al.
+Motivated by the limitations of single-judge evaluation, several works have explored multi-agent debate as a mechanism for approximating collective human judgment Hong et al.
 (2024).
-ChatEval was a pioneering framework in this area, introducing the idea of a “referee team” of LLM agents with diverse personas that debate to reach a consensus Chan et al.
+ChatEval Chan et al.
+(2024) introduced the idea of a referee team composed of LLM agents with distinct personas who debate before reaching a final verdict, demonstrating improved correlation with human judgments.
+PRD (Peer Rank and Discussion) Li et al.
+(2024) further refined this approach by emphasizing peer-review-style discussion to mitigate self-enhancement and positional biases Wang et al.
 (2023).
-This work established the potential of multi-agent debate to improve correlation with human judgments.
-PRD (Peer Rank and Discussion) further refined this by focusing on peer-review mechanisms to specifically target and reduce self-enhancement and positional biases Li et al.
-(2024); Wang et al.
+More recent work has expanded this line of research.
+KIEval (Yu et al., 2024) addresses data contamination through multi-round interactive dialogues grounded in domain knowledge, enabling evaluation of genuine model understanding over memorization.
+M-MAD (Feng et al., 2025) decomposes machine translation evaluation into four orthogonal error dimensions and uses multi-agent debate within each dimension to refine error detection and severity assessment, improving segment-level accuracy.
+In parallel, Zhang et al.
+(2025) demonstrate through comprehensive benchmarking that homogeneous multi-agent debate rarely outperforms simple baselines despite higher inference costs, arguing that model heterogeneity—not unconstrained deliberation—is essential for effective collaborative reasoning.
+D3 builds on these insights while addressing their practical limitations.
+Like ChatEval and PRD, it leverages role specialization and debate to reduce bias; however, it introduces a courtroom-inspired structure with explicitly defined Advocate, Judge, and Juror roles.
+Crucially, D3 differs from prior debate frameworks by explicitly managing the cost–accuracy trade-off through dual evaluation protocols and a budgeted stopping rule.
+This design directly responds to recent critiques of debate-based evaluation, enabling controlled deliberation rather than assuming that more agents or longer discussions are always beneficial.
+6.3 Specialized Evaluator Models
+An alternative to prompt-based or debate-driven evaluation is to train specialized models that act as judges Christiano et al.
+(2023); Ziegler et al.
+(2020); Bai et al.
 (2022).
-D3 builds directly on the insights from these systems.
-It incorporates the persona-driven approach of ChatEval and the bias-mitigation goals of PRD but introduces a more formalized, courtroom-inspired structure with distinct Advocate, Judge, and Juror roles.
-More importantly, D3’s primary novel contributions the explicit management of the cost-accuracy trade-off via its dual protocols and budgeted stopping rule, and its extensive, multi-benchmark empirical validation address key practical and scientific gaps left by these earlier frameworks.
-6.3 Specialized Evaluator Models
-An alternative to dynamic, prompt-based multi-agent systems is the development of specialized LLMs that are fine-tuned specifically for the task of evaluation Bai et al.
-(2022); Christiano et al.
-(2023); Ziegler et al.
+PandaLM exemplifies this approach as an open-source evaluator trained on human preference data to produce consistent and reproducible judgments Wang et al.
 (2024).
-PandaLM is a prominent open-source example, a model trained on human preference data to act as a reliable and reproducible judge Wang et al.
+Similarly, Prometheus aims to replicate GPT-4-level evaluation when conditioned on explicit evaluation criteria Kim et al.
 (2024).
-Similarly, Prometheus is another open-source LLM designed to replicate GPT-4’s evaluation capabilities when provided with appropriate criteria Kim et al.
+Specialized evaluators offer efficiency and stability advantages but are inherently limited by their training distribution and require retraining to adapt to new tasks or criteria.
+Recent analyses in 2025 have further highlighted their vulnerability to preference leakage and overfitting to benchmark-specific annotation styles Li et al.
+(2025); Marioriyad et al.
 (2025).
-These models offer the advantage of being smaller, faster, and potentially more consistent for a specific evaluation task.
-However, they may lack the flexibility of prompt-based systems to adapt to novel tasks or evaluation criteria without retraining.
-In our experiments, PandaLM serves as a strong non-debate baseline, allowing us to demonstrate that the dynamic, deliberative process of D3 can outperform even a highly specialized, fine-tuned model.
+In our experiments, PandaLM serves as a strong non-debate baseline, allowing us to isolate the contribution of D3’s deliberative process and show that structured, budget-aware debate can outperform even a highly optimized evaluator model.
 6.4 Automated Evaluation Benchmarks
-The empirical validation in this paper relies on the critical work done by the community in developing high-quality benchmarks Liang et al.
+This work relies on the substantial community effort devoted to building robust evaluation benchmarks Liang et al.
 (2023).
-MT-Bench provides a standardized testbed for general conversational abilities Zheng et al.
+MT-Bench provides a standardized benchmark for general conversational ability Zheng et al.
+(2023), while AlignBench offers a comprehensive multidimensional evaluation of alignment in Chinese, which we adapt for our study Liu et al.
 (2024).
-AlignBench fills a crucial gap by offering a comprehensive, multi-dimensional benchmark for evaluating alignment in Chinese, which we adapt for our study Liu et al.
+AUTO-J Li et al.
+(2023) introduces large-scale evaluation across 58 real-world scenarios using GPT-4 judgments, enabling broad empirical coverage.
+These benchmarks build on earlier work in automated metrics such as BLEU and ROUGE Papineni et al.
+(2002); Lin (2004), as well as long-standing critiques of their correlation with human judgment Callison-Burch et al.
+(2006); Mathur et al.
 (2020).
-Finally, AUTO-J introduces a large-scale dataset covering 58 real-world scenarios, with judgments provided by GPT-4, enabling evaluation across a wide spectrum of practical applications.
-Traditional metrics like BLEU Papineni et al.
-(2002), ROUGE Lin (2004), and their critiques Callison-Burch et al.
-(2006); Mathur et al.
-(2020) have informed the development of these modern benchmarks, while human evaluation standards Howcroft et al.
+Human evaluation protocols Howcroft et al.
 (2020); Celikyilmaz et al.
-(2021) and reading comprehension benchmarks like SQuAD Rajpurkar et al.
-(2016) have shaped evaluation methodology.
-By validating D3 across these diverse and challenging benchmarks, this work ensures that its demonstrated performance is robust and generalizable, supporting the broader goal of building general language assistants aligned with human values Askell et al.
+(2021) and task-specific benchmarks such as SQuAD Rajpurkar et al.
+(2016) have further shaped modern evaluation methodology.
+By validating D3 across multiple benchmarks with differing assumptions and failure modes, we aim to demonstrate that its improvements are robust rather than benchmark-specific, contributing to the broader goal of reliable and scalable alignment evaluation Askell et al.
 (2021).
-7 Discussion
-Our results show that structured, multi-agent frameworks such as D3 provide a more reliable and interpretable alternative to single-judge evaluations.
-By explicitly modeling the cost–accuracy frontier, D3 reframes evaluation as a trade-off rather than a single metric, helping practitioners select appropriate protocols for different development stages.
-This work thus highlights the need to treat cost as a core dimension in evaluation research.
-Beyond improving accuracy, future work should aim to make high-fidelity evaluations more accessible through distillation, automated role generation, and more efficient coordination mechanisms among agents.
 7 Conclusion
-In this paper, we introduced Debate, Deliberate, Decide (D3), a multi-agent framework designed to address critical gaps in LLM evaluation.
-Faced with the biases of single-LLM judges and the scalability limits of human annotation, D3 provides a structured, adversarial process that leverages role-specialized agents to produce more reliable assessments.
-Through a comprehensive empirical study, we demonstrated that D3 significantly outperforms strong baselines in agreement with human judgments.
-Crucially, our work moves beyond accuracy alone, presenting a systematic methodology for auditing biases and introducing the cost-accuracy frontier as a vital tool.
-By delivering a framework that is empirically robust, systematically audited, and practically cost-aware, D3 represents a significant step towards the scalable, reliable, and trustworthy evaluation of AI systems.
+D3 addresses critical gaps in LLM evaluation through structured, cost-aware multi-agent debate.
+Across MT-Bench, AlignBench, and AUTO-J, D3 outperforms baselines in accuracy, positional consistency, and self-enhancement robustness.
+D3-MORE provides efficiency comparable to ChatEval while achieving +8.3% higher accuracy than PRD.
+D3-SAMRE with budgeted stopping achieves highest accuracy at 4.65× single-judge cost, with 58% of debates converging by round 2.
+This cost-accuracy frontier enables practitioners to select protocols matching their constraints.
+By combining role specialization, systematic bias auditing, and explicit cost-awareness, D3 advances reliable and scalable LLM evaluation.
+Future work could explore automated role generation and distillation methods to further democratize access to high-fidelity evaluation.
 8 Limitations
@@
 Computational Cost.
-Although D3 is designed to be cost-aware, it remains substantially more expensive than single-judge evaluation.
+Although D3 is designed to be cost-aware, it remains more expensive than single-judge evaluation.
 The D3-MORE protocol requires roughly four times the tokens of a single-judge setup, and D3-SAMRE can consume even more.
@@
 Persona Design and Fairness.
-The use of diverse juror personas, while empirically beneficial, risks reinforcing social or cultural stereotypes.
-Responsible persona construction demands careful auditing to ensure diversity does not compromise fairness or neutrality.
-Developing principled, automated methods for persona generation remains an open challenge in human–AI interaction and responsible AI research.
+Our diverse juror personas provide empirical benefits (Section 5.1) but require responsible fairness management.
+Personas are role-based (ethics professor, business owner, social worker, environmental activist, tech entrepreneur) rather than demographic, reducing stereotype risks.
+All jurors use identical backbone LLM; personas only influence instructional framing.
+Section 5.1 validates robustness through controlled ablations: persona effects are stable, complementary ($89.5\%$ cross-persona agreement), and free of demographic bias (effect sizes <0.04<0.04).
+Open challenges remain: cross-cultural generalization in non-Western contexts, automated persona generation for scalability, and fairness audits across protected characteristics of response subjects.
+Future work should conduct targeted fairness studies to ensure D3 does not systematically advantage particular demographic groups.
 Dependence on Underlying Models.
@@
 Future work could explore distilling D3’s rationale-rich judgments into smaller, specialized evaluator models or introducing game-theoretic interactions among agents to enhance efficiency without sacrificing rigor.
+9 Acknowledgements
+We thank Amir Abdullah, Philip Quirke and Michael Lan for their invaluable feedback on the project.
+We also appreciate Antia Garcia Casal for her help with figures and for shaping the overall style of our visualizations.
 Appendix A Note on AI assistance
@@
 Your scores and explanation:
+Appendix G Persona Sensitivity Analysis: Full Ablation
+G.1 Systematic Persona Pool and Robustness Analysis
+We constructed a diverse pool of 50 personas spanning law, medicine, education, technology, ethics, business, social work, risk analysis, and compliance.
+We then conducted 10 independent experiments on MT-Bench (120 randomly selected questions), each using a different random 5-persona subset.
+All personas were instantiated with GPT-4-Turbo as the backbone; personas influenced only instructional prompts, not underlying model capabilities.
+Table 4 summarizes results:
+Table 4: Persona Robustness Analysis: Ablation Across Random 5-Persona Subsets (MT-Bench, 120 questions)
+Configuration Accuracy (%)
+Random 5-persona subsets (avg of 10) $85.9\pm 0.7$
+Curated 5-persona set $86.1$
+Generic jurors (no personas) $82.7\pm 1.1$
+(1) Persona Conditioning Effect:
+Persona-guided evaluation outperforms generic jurors by $3.2\%$ (85.9% vs.
+82.7%, $p<0.01$).
+The persona effect holds robustly across random subsets, indicating diversity, not specific persona identity, drives performance.
+(2) Robustness Across Subsets:
+Tight variance ($0.7\%$ std. dev.) indicates D3’s performance is not brittle to persona composition; the framework generalizes well to unseen persona combinations.
+(3) Curated vs. Random:
+Curated personas ($86.1\%$) achieve marginally higher accuracy than the random average ($85.9\%$, $0.2\%$). This validates that deliberate design adds value without harming generalization.
+G.2 Design Rationale: Persona Selection and Value Lenses
+Our five curated personas were selected to provide complementary professional perspectives and value lenses.
+Each persona is anchored in domain expertise rather than demographic attributes, reducing stereotype risks:
+Framework Evaluator Accuracy (%) Relative Gain
+Single Judge Llama-3-70B 68.3 baseline
+ChatEval Llama-3-70B 73.1 +4.8%
+PRD Llama-3-70B 71.5 +3.2%
+D3-MORE Llama-3-70B 73.9 +5.6%
+D3-MORE GPT-4-Turbo 85.1 +12.6%
+Table 5: MT-Bench accuracy using open-source and proprietary evaluators.
+  • Ethics Professor: Ethical principles, long-term societal impact
+  • Environmental Activist: Collective welfare, ecological responsibility
+  • Business Owner: Practical feasibility, roi, trade-offs
+  • Social Worker: Human-centered perspective, equity, vulnerable populations
+  • Tech Entrepreneur: Innovation, scalability, technological progress
+This composition ensures verdicts incorporate ethical grounding, distributional impact, economic viability, and technological feasibility.
+These roles are intentionally role-based rather than demographic; all jurors execute the same backbone model with identical base capabilities.
+G.3 Cross-Persona Agreement and Complementarity
+We computed pairwise vote agreement across the five curated personas.
+Average agreement is $89.5\%\pm 1.5\%$, with pairwise ranges $87{-}92\%$.
+High cross-persona agreement confirms personas are complementary rather than contradictory, reducing the risk of arbitrarily conflicting judgments while maintaining diverse perspectives.
+G.4 SAMRE Debate Progression Across Question Types
+The convergence behavior of iterative debate varies systematically across task complexity and reasoning demands.
+We analyze score gap trajectories across MT-Bench question categories to validate that the budgeted stopping rule terminates debates at their point of maximal signal without sacrificing verdict reliability.
+Coding tasks exhibit the largest discriminative gaps, with peaks reaching 20 points by round 2 and maintaining plateau stability through round 5.
+This reflects clear correctness boundaries in code evaluation.
+Reasoning and ethics questions show moderate, steady gaps (8–11 points) with gentle convergence, indicating these subjective domains benefit from iterative refinement but stabilize predictably.
+Writing, roleplay, and math tasks display tighter gaps (0–6 points) and earlier plateau behavior, suggesting these domains reach verdict saturation with fewer debate rounds.
+Across all categories, gap stabilization occurs by rounds 4–5, with the majority of verdicts determined by round 2.
+The final round markers (diamonds) per question type show that D3-SAMRE can terminate heterogeneously: early for low-variance tasks like math and writing, while extending slightly longer for nuanced reasoning and ethics evaluations.
+This task-sensitive convergence supports the empirical guidance in Section 5.5: use SAMRE for high-stakes and ethically complex evaluations, while MORE suffices for objective or well-separated task types.
+Figure 5: Score gap trajectories across six MT-Bench question categories reveal differential convergence patterns in iterative adversarial debate.
+Coding tasks exhibit the largest gaps (peak 20 points) with stable plateaus, indicating clear discriminability.
+Reasoning and ethics tasks show moderate, steady gaps (8–11 points).
+Writing, roleplay, and math questions display tighter gaps (0–6 points) with earlier plateau behavior.
+Diamond markers indicate final verdict round per category.
+Appendix H Open-Source Evaluator Results
+To test whether D3’s performance gains depend on evaluator model quality, we evaluate all frameworks using Llama-3-70B Llama (2024) as the evaluator on 100 MT-Bench Zheng et al.
+(2023) questions.
+Table 5 reports accuracy and relative gains over a single-judge baseline.
+These results show that D3’s improvements persist when using a fully open-source evaluator, indicating that the gains stem from the evaluation architecture rather than reliance on proprietary judge models.